The Griggs Corporation has credit sales of $1,075,900 Total assets turnover Cash to total assets Accounts receivable turnower 2.90 times 1.30% Inventory turnover Current ratio Debt to total assets 10 times 4 times 1.83 times 45% Using the above ratios, complete the balance sheet. (Round your answers to the nearest whole number.) GRIGGS CORPORATION Balance Sheet Assets Liabilities and Stockholders Equity Cash Accounts receivable Inventory Current debt Long-term debt Total debt Total current assets Fixed assets Equity Total assets Total debt and stockholders equity
Answer #1
As cost of goods sold is not given, therefore sales revenue will taken in the formula of inventory turnover ratio.
Net sales: 1075900
Total assets turnover =2.90 times
Total assets turnover = Net sales/ Total assets
2.90 = 1075900 / Total assets
Total assets = 371000
Cash to Total assets = 1.30%
Cash = 371000 *1.30% = 4823
Accounts receivable turnover = 10 times
Accounts receivable turnover = Net sales / Accounts receivable
10 = 1075900 / Accounts receivable
Accounts receivable = 107590
Inventory turnover ratio =14 times
Inventory Turnover ratio = Net sales/ Inventory
14 = 1075900 / Inventory
Inventory =76850
Total current assets = Cash+Accounts receivable + Inventory
4823+107590+76850 = 189263
Total assets = Total current assets+ Fixed assets
371000 = 189263 - Fixed assets
Fixed assets = 371000-189263 =181737
Current ratio =1.88 times
Current ratio = Current assets / Current liabilties
1.88 = 189263 / Current liabilities
Current liabilities = 100672
Debt to t otal assets =45%
Debt to total assets = Debt / Total assets
Debts = 371000*45% = 166950
Long term debt = Total debts - Current liabilities = 166950-100672 = 66278
Total liabilities and Stockholder equity = Total assets = 371000
Therefore
Stockholder's equity = 371000 - Total debts =371000-166950 = 204050
Balance Sheet:
Assets Liabilities and Stockholder's equity
Cash 4823 Current liabilities 100672
Accounts receivable 107950 Long term debts 66278
Inventory 76850 Total liabilities 166950
Total current assets 189263
Fixed assets 181737 Equity 204050
Total Assets 371000 Total debt and Stockholder's equity 371000